# Finding the unusual red giant remnants of cataclysmic variable mergers

### v1

@@ 4.5 Populating the horizontal branch with merger remnants
Neglecting mass loss during merger or through winds, $M_{\mathrm{MS}}$ sets the envelope mass of the merger remnant. In the preceding discussion, we have focused on varying $M_{\mathrm{WD}}$ and fixed $M_{\mathrm{MS}}$ so that the total mass of the merger remnant models is $M=M_{\mathrm{WD}}+M_{\mathrm{MS}}=0.80M_{\odot}$. However, in principle, $M_{\mathrm{MS}}$ could span a wide range of masses, including very small ones (if the hydrogen-rich component is a brown dwarf or if there is significant mass loss, e.g., Metzger et al. 2021), up to a few solar masses for mergers with intermediate-mass MS stars. In this section, we explore the behavior of merger remnants under varying $M_{\mathrm{MS}}$.
@@
    If $M_{\mathrm{MS}}$ is sufficiently small, merger remnants may only ignite helium with scant hydrogen envelopes (starting CHeB as subdwarf B-type stars; sdBs), or may fail to ignite helium altogether (fizzling out into He WDs). For reasons of scope, we do not investigate the He WD or sdB outcomes (but see the detailed modelling of Zhang et al. 2017 and Zhang et al. 2023). Alternatively, if the envelope mass drops below $M_{\mathrm{env}}\simeq 0.1M_{\odot}$ during CHeB but can still sustain hydrogen shell burning, the remnant can evolve onto the horizontal branch (Catelan, 2009).
@@ 5.1 Undermassive red clump stars may be merger remnants
Using asteroseismology, Li et al. (2022) discovered a population of undermassive CHeB stars with masses $\lesssim 0.8M_{\odot}$ (their “very low-mass” sample). Since single stars of these masses could not have evolved off of the MS in the age of the universe, these undermassive stars must have undergone non-standard evolution, such as stripping by a companion or binary assembly. Although Li et al. (2022) argue that these undermassive giants are the product of partial envelope stripping by close companions, most of these objects do not exhibit the expected radial velocity variability between APOGEE and Gaia (Kareem El-Badry, private communication), disfavoring this formation channel. Matteuzzi et al. (2023) identify several more such objects (referred to as “red horizontal branch” stars), further demonstrating their extremely strong mixed-mode coupling.
@@ 5.2 Zvrk: a possible post-dredge-up merger remnant?
    The star is highly lithium-rich ($\mathrm{A(Li)}>3$), and also slightly carbon-enhanced relative to typical first-ascent RGs, or indeed CHeB stars (e.g., Bufanda et al., 2023).
@@
    As demonstrated in Section 4.4, post-dredge-up remnants are expected to be carbon-rich, similar to Zvrk. While the core dredge-up event at the helium flash is unlikely to match the measured value of $\mathrm{A}(\mathrm{Li})$, lithium enrichment may occur at merger (see Section 4.4).
@@
    As discussed in Section 6.3, merger remnants are expected to be rapidly rotating.
It is unclear at present whether He WD–MS mergers can reproduce these effects in the correct combination to match observations. The mass $M\approx 1.2M_{\odot}$ inferred by scaling relations is larger than those of our fiducial models ($M=0.8M_{\odot}$), possibly requiring even stronger core dredge-ups which probably occur for larger values of $M_{\mathrm{WD}}$ than we can model. We point out that an extreme dredge-up event of this type probably modifies $\nu_{\mathrm{max}}$ (in a still-contested manner; Viani et al., 2017; Zhou et al., 2024) and, thus, the accuracy of scaling relation-based values of $M$ and $R$. The similarly behaving remnants of RG–He WD mergers Zhang & Jeffery (e.g., 2013) may also possibly reproduce the properties of Zvrk.
@@ 5.3 Other potential post-dredge-up merger remnants
Recently, a small number of highly carbon-deficient giants has been discovered and characterized as a distinct class (Bidelman & MacConnell, 1973; Bond, 2019; Maben et al., 2023). These objects are almost all in the CHeB phase, and many are also lithium-rich and overluminous compared to the usual red clump (Maben et al., 2023). However, the He WD–MS merger scenario predicts a carbon enrichment and thus fails to explain the abnormally low carbon abundances in these stars.
@@ 6.1 Progenitors and rates
While no CVs with He WD accretors have been discovered (to our knowledge), CV progenitors (known as pre-CVs) have been, some shown in Figure 13. As their name suggests, such systems are expected to eventually initiate (possibly unstable) mass transfer after a combination of magnetic braking and gravitational radiation tighten their orbits sufficiently.
@@
Based on MESA models of merger remnants (Zhang et al., 2017), Zhang et al. (2018) perform a population synthesis calculation to determine how much these sorts of merger remnants contribute to the population of single He WDs, which are not a natural outcome of isolated stellar evolution. Assuming a star-formation rate of $5M_{\odot}\,\mathrm{yr}^{-1}$, their calculation implies a Galactic formation rate of merger remnants $\sim 0.02\,\mathrm{yr}^{-1}$ (about half of which fizzle out into He WDs before starting helium burning), with a factor-of-a-few uncertainty when the Reimers wind parameter is varied (Reimers, 1975, 1977). If merger remnants continue to appear as RGs for $100T_{\mathrm{m}8}\,\mathrm{Myr}$, the Milky Way should contain $\sim 2T_{\mathrm{m}8}\times 10^{6}$ remnants at a given time. Put another way, for a Kroupa initial mass function (Kroupa, 2001), single low-mass RGs in the range $0.8M_{\odot}$ to $2.0M_{\odot}$ form at a rate $\sim 0.3M_{\odot}\,\mathrm{yr}^{-1}$. If isolated RGs have lifetimes $100T_{\mathrm{s}8}\,\mathrm{Myr}$, He WD–MS merger remnants should make up roughly $\sim 7\%\times(T_{\mathrm{m}8}/T_{\mathrm{s}8})$ of all low-mass RGs.
@@ 6.3 Rapid rotation
where $M=M_{\mathrm{MS}}+M_{\mathrm{WD}}$ (and we have assumed a circular orbit). Then, assuming that no mass is lost during the merger, the final envelope spin rate of the resulting RG remnant is
@@ 6.5 Broader progress in binary interaction asteroseismology
  • RG–He WD mergers likely produce unusual CHeB giants, which have been previously explored as a possible channel for producing certain classes of carbon stars (Izzard et al., 2007; Zhang & Jeffery, 2013; Zhang et al., 2020) or as a possible progenitor of the 8 UMi planetary system (Hon et al., 2023). The models of Zhang & Jeffery (2013) suggest that these merger remnants behave like overluminous CHeB stars, similar to those described in Section 4.1 for the He WD–MS scenario (compare their Figure 4 to our Figure 2).
  • CO WD–MS mergers are likely to result from consequential angular momentum loss, particularly for lower-mass CO WDs. While existing modelling literature typically focuses on progenitor systems’ nova eruptions (Iben & Tutukov, 1996; Shara et al., 2010; Kato et al., 2017) or their role in producing Type Ia supernovae (Kovetz & Prialnik, 1994; Cassisi et al., 1998; Newsham et al., 2013; Hillman et al., 2016; Wang, 2018), mergers of such remnants should become unusual asymptotic giant branch stars with highly degenerate cores (Cassisi et al., 1998; Piersanti et al., 2000; Wolf et al., 2013), with possibly observable consequences.
  • CO WD–He WD mergers are the likely progenitors of R Coronae Borealis stars (R Cor Bor stars), which are yellow supergiants consisting of a carbon–oxygen core surrounded by a helium envelope inflated by shell burning. R Cor Bor stars are known to pulsate at periods between $40$ and $100\,\mathrm{d}$ (Lawson & Kilkenny, 1996; Karambelkar et al., 2021), making asteroseismology a promising tool for probing their internal structures (Wong & Bildsten, 2023).
  • Last year, Bellinger et al. (2023) identified asteroseismology as a tool for testing the post-MS merger channel for producing blue supergiants, finding that the g-mode period spacing $\Delta\Pi$ constrains their formation channels. In a parallel observational study using TESS, Ma et al. (2023) discovered peculiar but universal low-frequency ($f\lesssim 2\,\mathrm{d}^{-1}$) photometric power excess, although the physical interpretation remains unclear.
@@ 7 Summary
We thank Kareem El-Badry, Joel Ong, Marc Hon, and Yaguang Li for their insightful comments and thorough reading of the manuscript, as well as Lars Bildsten for useful discussions. We also acknowledge Masao Takata, Andrew Casey, and Tim Bedding for their helpful remarks. N.Z.R. acknowledges support from the National Science Foundation Graduate Research Fellowship under Grant No. DGE‐1745301.
Appendix A Dependence on the cooling age of the white dwarf
@@
    A merger remnant with $\log(L_{\mathrm{WD}}/L_{\odot})=-4.0$ and a lower mass $M_{\mathrm{WD}}=0.27M_{\odot}$.
@@
Appendix B Thermal timescale of the core
@@
$\vec{q}=-\alpha_{\mathrm{cond}}\rho T\nabla T$ (B1)
@@
$\alpha_{\mathrm{cond}}\approx 2.44\times 10^{3}AZ^{-2}\,\mathrm{cm}^{4}\,\mathrm{s}^{-3}\,\mathrm{K}^{-2}\mathrm{,}$ (B2)
@@
$c_{v}\approx\frac{\varepsilon}{T}\approx\frac{3\rho k_{B}}{2Am_{p}}$ (B3)
@@
$\frac{\partial\varepsilon}{\partial t}=-\nabla\cdot\vec{q}$ (B4)
Substituting Equations B1 and B3 into Equation B4 yields the following nonlinear heat diffusion equation:
$\frac{\partial T}{\partial t}=\frac{\kappa}{r^{2}}\frac{\partial}{\partial r}\left(r^{2}T\frac{\partial T}{\partial r}\right)\mathrm{,}$ (B5)
cf. Mestel (1952). In writing Equation B5, we have assumed spherical symmetry, and that $\rho$ does not vary with temperature. The diffusion coefficient $\kappa$ is given by
$\kappa=\frac{2Am_{p}\alpha_{\mathrm{cond}}}{3k_{B}}$ (B6)
At early times, the core is essentially isothermal at a low temperature $T_{0}\ll T_{\mathrm{shell}}\simeq 3\times 10^{7}\,\mathrm{K}$. Thermal contact with the hot hydrogen-burning shell at the outer boundary sets up a temperature gradient $\partial T/\partial r\sim T_{\mathrm{shell}}/R_{c}$. The natural length scale of the problem is $R_{c}$. By Equation B5, the core therefore thermalizes on a timescale $\tau_{\mathrm{th,core}}$ given roughly by
$\frac{1}{\tau_{\mathrm{th,core}}}\sim\frac{\kappa T_{\mathrm{shell}}}{R_{c}^{2}}$ (B7)
@@
$\tau_{\mathrm{th,core}}\sim\frac{3k_{B}}{2Am_{p}\alpha_{\mathrm{cond}}}\frac{R_{c}^{2}}{T_{\mathrm{shell}}}\approx 80\,\mathrm{Myr}\times\left(\frac{R_{c}}{0.035R_{\odot}}\right)^{2}\left(\frac{T_{\mathrm{shell}}}{3\times 10^{7}\,\mathrm{K}}\right)^{-1}\left(\frac{Z}{2}\right)^{2}\left(\frac{A}{4}\right)^{-2}$ (B8)
A similar timescale to $\tau_{\mathrm{th,core}}$ has been previously derived by Shen et al. (2009). The scaling relation in Equation B8 is normalized to a typical He WD in thermal contact with a typical hydrogen-burning shell. This timescale is comparable to the length of the remnant’s RGB phase—the longest RGB phase leading to helium ignition in our merger remnant models lasts $\simeq 180\,\mathrm{Myr}$ (for a He WD mass $M_{\mathrm{WD}}=0.20M_{\odot}$). Heat conduction thus cannot completely destroy the low-entropy state of the core quickly enough to erase its long-term effects on the appearance and evolution of the remnant, especially for higher values of $M_{\mathrm{WD}}$.
Appendix C Brunt–Väisälä frequency in degenerate helium cores
@@
$N^{2}=N_{0}^{2}\frac{\chi_{T}}{\chi_{\rho}}\left(\nabla_{\mathrm{ad}}-\nabla\right)$ (C1)
@@
$\chi_{T}=\left(\frac{\partial\ln p}{\partial\ln T}\right)_{\rho}$ (C2a)
$\chi_{\rho}=\left(\frac{\partial\ln p}{\partial\ln\rho}\right)_{T}$ (C2b)
$\nabla_{\mathrm{ad}}=\left(\frac{\mathrm{d}\ln T}{\mathrm{d}\ln p}\right)_{\mathrm{ad}}$ (C2c)
$\nabla=\frac{\mathrm{d}\ln T}{\mathrm{d}\ln p}$ (C2d)
@@
$p\approx\frac{2}{5}\frac{Z\rho E_{F}}{Am_{p}}$ (C5a)
$\left(\frac{\partial p}{\partial T}\right)_{\rho}\approx\frac{\rho k_{B}}{Am_{p}}$ (C5b)
@@
$\chi_{T}=\frac{T}{p}\left(\frac{\partial p}{\partial T}\right)_{\rho}=\frac{5}{2Z}\frac{k_{B}T}{E_{F}}\mathrm{.}$ (C6)
@@
$N^{2}\approx N_{0}^{2}\frac{k_{B}T}{ZE_{F}}\left(1-\frac{5}{2}\nabla\right)\mathrm{.}$ (C7)

### v2

@@ 4.5 Populating the horizontal branch with merger remnants
Neglecting mass loss during merger or through winds, $M_{\mathrm{MS}}$ sets the envelope mass of the merger remnant. In the preceding discussion, we have focused on varying $M_{\mathrm{WD}}$ and fixed $M_{\mathrm{MS}}$ so that the total mass of the merger remnant models is $M_{\rm tot}=M_{\mathrm{WD}}+M_{\mathrm{MS}}=0.80M_{\odot}$. However, in principle, $M_{\mathrm{MS}}$ could span a wide range of masses, including very small ones (if the hydrogen-rich component is a brown dwarf or if there is significant mass loss, e.g., Metzger et al. 2021), up to a few solar masses for mergers with intermediate-mass MS stars. In this section, we explore the behavior of merger remnants under varying $M_{\mathrm{MS}}$.
@@
    If $M_{\mathrm{MS}}$ is sufficiently small, merger remnants may only ignite helium with scant hydrogen envelopes (starting CHeB as subdwarf B-type stars; sdBs), or may fail to ignite helium altogether (fizzling out into He WDs). For reasons of scope, we do not investigate the He WD or sdB outcomes (but see the detailed modeling of Zhang et al. 2017 and Zhang et al. 2023). Alternatively, if the envelope mass drops below $M_{\mathrm{env}}\simeq 0.1M_{\odot}$ during CHeB but can still sustain hydrogen shell burning, the remnant can evolve onto the horizontal branch (Catelan, 2009).
@@ 5.1 Undermassive red clump stars may be merger remnants
Using asteroseismology, Li et al. (2022) discovered a population of undermassive CHeB stars with masses $\lesssim 0.8M_{\odot}$ (their “very low-mass” sample). Since single stars of these masses could not have evolved off of the MS in the age of the universe, these undermassive stars must have undergone non-standard evolution, such as stripping by a companion or binary assembly. Although Li et al. (2022) argue that these undermassive giants are the product of partial envelope stripping by close companions, most of these objects do not exhibit the expected radial velocity variability between APOGEE and Gaia (Kareem El-Badry, private communication), disfavoring this formation channel. Matteuzzi et al. (2023) identify several more such objects (referred to as “red horizontal branch” stars), further demonstrating their extremely strong mixed-mode coupling. Figure 5 shows $8$ members of the very low-mass sample of Li et al. (2022) for which Vrard et al. (2016) reports g-mode period spacings. Despite their small masses, these very low-mass CHeB stars have typical values of $\Delta\Pi$ and, therefore, likely possess similar core structures to those of normal CHeB stars.
@@ 5.2 Zvrk: a possible post-dredge-up merger remnant?
    The star is highly lithium-rich ($\mathrm{A(Li)}>3$), and also has a somewhat high [C/N] ratio relative to typical first-ascent RGs of the same mass, or indeed CHeB stars (e.g., Bufanda et al., 2023).
@@
    As demonstrated in Section 4.4, post-dredge-up remnants are expected to be carbon-rich, similar to Zvrk. While the core dredge-up event at the helium flash is unlikely to match the measured value of $\mathrm{A}(\mathrm{Li})$, lithium enrichment may occur at merger (Kamiński et al., 2023, as suggested by observations of luminous red novae). Although some of our merger remnant models become very lithium-rich soon after merger (as we describe in Section 4.4), future work should address whether this enhancement persists under more careful modeling.
@@
    As discussed in Section 6.3, merger remnants are likely rapidly rotating.
It is unclear at present whether He WD–MS mergers can reproduce these effects in the correct combination to match observations. The mass $M\approx 1.2M_{\odot}$ inferred by scaling relations is larger than those of our fiducial models ($M=0.8M_{\odot}$), possibly requiring even stronger core dredge-ups which probably occur for larger values of $M_{\mathrm{WD}}$ than we can model. We point out that an extreme dredge-up event of this type probably modifies $\nu_{\mathrm{max}}$ (in a still-contested manner; Viani et al., 2017; Zhou et al., 2024) and, thus, the accuracy of scaling relation-based values of $M$ and $R$. The similarly behaving remnants of RG–He WD mergers (e.g., Zhang & Jeffery, 2013) may also possibly reproduce the properties of Zvrk.
@@ 5.3 Other potential post-dredge-up merger remnants
Recently, a small number of highly carbon-deficient red giants has been discovered and characterized as a distinct class with a few common properties (Bidelman & MacConnell, 1973; Bond, 2019; Maben et al., 2023). These objects are almost all in the CHeB phase (as implied by their g-mode period spacings), and many are also lithium-rich and overluminous compared to the usual red clump (Maben et al., 2023). However, the He WD–MS merger scenario predicts a carbon enrichment and thus fails to explain the abnormally low carbon abundances in these stars.
@@ 6.1 Progenitors and rates
While very few CVs with He WD accretors have been discovered (to our knowledge), CV progenitors (known as pre-CVs) have been, some shown in Figure 14. As their name suggests, such systems are expected to eventually initiate (possibly unstable) mass transfer after a combination of magnetic braking and gravitational radiation tighten their orbits sufficiently.
@@
Based on MESA models of merger remnants (Zhang et al., 2017), Zhang et al. (2018) perform a population synthesis calculation to determine how much these sorts of merger remnants contribute to the population of single He WDs, which are not a natural outcome of isolated stellar evolution (see also Zorotovic & Schreiber, 2017). Assuming a star-formation rate of $5M_{\odot}\,\mathrm{yr}^{-1}$, their calculation implies a Galactic formation rate of merger remnants $\sim 0.02\,\mathrm{yr}^{-1}$ (about half of which fizzle out into He WDs before starting helium burning), with a factor-of-a-few uncertainty when the Reimers wind parameter is varied (Reimers, 1975, 1977). If merger remnants continue to appear as RGs for $100T_{\mathrm{m}8}\,\mathrm{Myr}$, the Milky Way should contain $\sim 2T_{\mathrm{m}8}\times 10^{6}$ remnants at a given time. Put another way, for a Kroupa initial mass function (Kroupa, 2001), single low-mass RGs in the range $0.8M_{\odot}$ to $2.0M_{\odot}$ form at a rate $\sim 0.3M_{\odot}\,\mathrm{yr}^{-1}$. If isolated RGs have lifetimes $100T_{\mathrm{s}8}\,\mathrm{Myr}$, He WD–MS merger remnants should make up roughly $\sim 7\%\times(T_{\mathrm{m}8}/T_{\mathrm{s}8})$ of all low-mass RGs.
The present work has focused on the case where the compact component is a He WD. However, in principle, lower-mass CO WDs may also participate in mergers with MS stars for the same reasons. In such cases, the MS component is expected to disrupt around and accrete onto the carbon–oxygen core and initiate shell burning, and the merger remnant is therefore likely to resemble an asymptotic giant branch star with an unusually cold core. We defer a detailed investigation of this interesting possibility to a future work.
@@ 6.3 Rapid rotation
While most RG envelopes rotate very slowly, spectroscopic (Carlberg et al., 2011) and photometric (Ceillier et al., 2017) studies have indicated that $\sim 2\%$ of RGs are rapidly rotating ($v\sin i\gtrsim 10\,\mathrm{km}\,\mathrm{s}^{-1}$, or $v/v_{\mathrm{crit}}\gtrsim 7\%$ for typical values $M\sim M_{\odot}$ and $R\sim 10R_{\odot}$), likely as a result of stellar interactions such as mergers. The He WD–MS merger remnants discussed in this work may also rotate rapidly enough for rotation measurements to serve as an orthogonal diagnosis for their binary origin.
@@
where $M_{\mathrm{tot}}=M_{\mathrm{MS}}+M_{\mathrm{WD}}$ (and we have assumed a circular orbit). Then, assuming that no mass is lost during the merger, the final envelope spin rate of the resulting RG remnant is
@@
Throughout this work, we have neglected rotation-induced mixing processes (Zahn, 1994; Talon et al., 1997; Mathis & Zahn, 2004; Zahn, 2010; Park et al., 2020) which may be important in a rapidly rotating merger remnant. Our predictions for the amount of helium and other species mixed into the envelope and brought to the surface during a core dredge-up event should thus be considered lower limits.
@@ 6.5 Broader progress in binary interaction asteroseismology
  • RG–He WD mergers likely produce unusual CHeB giants, which have been previously explored as a possible channel for producing certain classes of carbon stars (Izzard et al., 2007; Zhang & Jeffery, 2013; Zhang et al., 2020) and as possible progenitors of the 8 UMi planetary system (Hon et al., 2023) and CK Vulpeculae, a historical transient observed in the year 1670 which is now a bipolar nebula (Tylenda et al., 2024). The models of Zhang & Jeffery (2013) suggest that these merger remnants behave like overluminous CHeB stars, similar to those described in Section 4.1 for the He WD–MS scenario (compare their Figure 4 to our Figure 3).
  • CO WD–MS mergers are likely to result from consequential angular momentum loss, particularly for lower-mass CO WDs. While existing modeling literature typically focuses on progenitor systems’ nova eruptions (Iben & Tutukov, 1996; Shara et al., 2010; Kato et al., 2017) or their role in producing Type Ia supernovae (Kovetz & Prialnik, 1994; Cassisi et al., 1998; Newsham et al., 2013; Hillman et al., 2016; Wang, 2018), mergers of such remnants should become unusual asymptotic giant branch stars with highly degenerate cores (Cassisi et al., 1998; Piersanti et al., 2000; Wolf et al., 2013), with possibly observable consequences.
  • CO WD–He WD mergers are the likely progenitors of R Coronae Borealis stars (R Cor Bor stars), which are yellow supergiants consisting of a carbon–oxygen core surrounded by a helium envelope inflated by shell burning (Clayton et al., 2007; Menon et al., 2013). R Cor Bor stars are known to pulsate at periods between $40$ and $100\,\mathrm{d}$ (Lawson & Kilkenny, 1996; Karambelkar et al., 2021), making asteroseismology a promising tool for probing their internal structures (Wong & Bildsten, 2023).
  • Last year, Bellinger et al. (2023) identified asteroseismology as a tool for testing the post-MS merger channel for producing blue supergiants, finding that the g-mode period spacing $\Delta\Pi$ constrains their formation channels. In a parallel observational study using TESS, Ma et al. (2023) discovered a peculiar but universal low-frequency ($f\lesssim 2\,\mathrm{d}^{-1}$) photometric power excess, although the physical nature of these oscillations remains unclear, and the authors were unable to observe individual modes or measure $\Delta\Pi$.
@@ 7 Summary
We thank Kareem El-Badry, Joel Ong, Marc Hon, and Yaguang Li for their insightful comments and thorough reading of the manuscript, as well as Lars Bildsten for useful discussions. We also acknowledge Masao Takata, Andrew Casey, Tim Bedding, and Ken Shen for their helpful remarks, as well as the anonymous referee whose report increased the clarity of the final manuscript. N.Z.R. acknowledges support from the National Science Foundation Graduate Research Fellowship under Grant No. DGE‐1745301. All stellar evolution calculations were performed on the Wheeler cluster at Caltech, which was supported by the Sherman Fairchild Foundation and by Caltech.
Appendix A A. Dependence on the cooling age of the white dwarf
@@
    A merger remnant with $\log(L_{\mathrm{WD}}/L_{\odot})=-4.0$ and a lower-mass $M_{\mathrm{WD}}=0.27M_{\odot}$.
@@
Appendix B B. Thermal timescale of the core
@@
$\vec{q}=-\alpha_{\mathrm{cond}}\rho T\nabla T$ (14)
@@
$\alpha_{\mathrm{cond}}\approx 2.44\times 10^{3}AZ^{-2}\,\mathrm{cm}^{4}\,\mathrm{s}^{-3}\,\mathrm{K}^{-2}\mathrm{,}$ (15)
@@
$c_{v}\approx\frac{\varepsilon}{T}\approx\frac{3\rho k_{B}}{2Am_{p}}$ (16)
@@
$\frac{\partial\varepsilon}{\partial t}=-\nabla\cdot\vec{q}$ (17)
Substituting Equations 14 and 16 into Equation 17 yields the following nonlinear heat diffusion equation:
$\frac{\partial T}{\partial t}=\frac{\kappa}{r^{2}}\frac{\partial}{\partial r}\left(r^{2}T\frac{\partial T}{\partial r}\right)\mathrm{,}$ (18)
cf. Mestel (1952). In writing Equation 18, we have assumed spherical symmetry, and that $\rho$ does not vary with temperature. The diffusion coefficient $\kappa$ is given by
$\kappa=\frac{2Am_{p}\alpha_{\mathrm{cond}}}{3k_{B}}$ (19)
At early times, the core is essentially isothermal at a low temperature $T_{0}\ll T_{\mathrm{shell}}\simeq 3\times 10^{7}\,\mathrm{K}$. Thermal contact with the hot hydrogen-burning shell at the outer boundary sets up a temperature gradient $\partial T/\partial r\sim T_{\mathrm{shell}}/R_{c}$. The natural length scale of the problem is $R_{c}$. By Equation 18, the core therefore thermalizes on a timescale $\tau_{\mathrm{th,core}}$ given roughly by
$\frac{1}{\tau_{\mathrm{th,core}}}\sim\frac{\kappa T_{\mathrm{shell}}}{R_{c}^{2}}$ (20)
@@
$\tau_{\mathrm{th,core}}\sim\frac{3k_{B}}{2Am_{p}\alpha_{\mathrm{cond}}}\frac{R_{c}^{2}}{T_{\mathrm{shell}}}\approx 80\,\mathrm{Myr}\times\left(\frac{R_{c}}{0.035R_{\odot}}\right)^{2}\left(\frac{T_{\mathrm{shell}}}{3\times 10^{7}\,\mathrm{K}}\right)^{-1}\left(\frac{Z}{2}\right)^{2}\left(\frac{A}{4}\right)^{-2}$ (21)
A similar timescale to $\tau_{\mathrm{th,core}}$ has been previously derived by Shen et al. (2009). The scaling relation in Equation 21 is normalized to a typical He WD in thermal contact with a typical hydrogen-burning shell. This timescale is comparable to the length of the remnant’s RGB phase—the longest RGB phase leading to helium ignition in our merger remnant models lasts $\simeq 180\,\mathrm{Myr}$ (for a He WD mass $M_{\mathrm{WD}}=0.20M_{\odot}$). Heat conduction thus cannot completely destroy the low-entropy state of the core quickly enough to erase its long-term effects on the appearance and evolution of the remnant, especially for higher values of $M_{\mathrm{WD}}$.
Appendix C C. Brunt–Väisälä frequency in degenerate helium cores
@@
$N^{2}=N_{0}^{2}\frac{\chi_{T}}{\chi_{\rho}}\left(\nabla_{\mathrm{ad}}-\nabla\right)$ (22)
@@
$\chi_{T}=\left(\frac{\partial\ln p}{\partial\ln T}\right)_{\rho}$ (23a)
$\chi_{\rho}=\left(\frac{\partial\ln p}{\partial\ln\rho}\right)_{T}$ (23b)
$\nabla_{\mathrm{ad}}=\left(\frac{\mathrm{d}\ln T}{\mathrm{d}\ln p}\right)_{\mathrm{ad}}$ (23c)
$\nabla=\frac{\mathrm{d}\ln T}{\mathrm{d}\ln p}$ (23d)
@@
$p\approx\frac{2}{5}\frac{Z\rho E_{F}}{Am_{p}}$ (26a)
$\left(\frac{\partial p}{\partial T}\right)_{\rho}\approx\frac{\rho k_{B}}{Am_{p}}$ (26b)
@@
$\chi_{T}=\frac{T}{p}\left(\frac{\partial p}{\partial T}\right)_{\rho}=\frac{5}{2Z}\frac{k_{B}T}{E_{F}}\mathrm{.}$ (27)
@@
$N^{2}\approx N_{0}^{2}\frac{k_{B}T}{ZE_{F}}\left(1-\frac{5}{2}\nabla\right)\mathrm{.}$ (28)
Appendix D D. Post-helium flash composition profiles
Figure 16: Composition profiles for the non-merged $M=0.80M_{\odot}$ and selected merger remnant models at the beginning of CHeB (a few megayear after the helium flash).
Figure 16 compares the composition profiles for various merger remnants at the beginning of CHeB, soon after the helium flash. As explained in Section 4, merger remnants involving more massive He WDs (higher $M_{\mathrm{WD}}$) convert a larger fraction of their core helium into carbon during the helium flash. Moreover, our merger remnant models with $M_{\mathrm{WD}}\geq 0.27M_{\odot}$ (e.g., the bottom two panels of Figure 16) undergo core dredge-up events which mix some of the helium core into the envelope. Such merger remnants therefore begin CHeB with lower core masses and possess envelopes which are dramatically enhanced in species such as helium and carbon (see Section 4.4).